# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

### v3

@@ 3.3. ReDoS Definitions
Table 3. Summary of ReDoS vulnerability definitions in reviewed studies. The literature uses inconsistent criteria, with over half (19/29) defining ReDoS based on slowdown.
Paper Slowdown Number of Instructions Heuristics Su et al. (Su et al., 2024) ¿10000 Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) 10s, 100s Liu et al. (Liu et al., 2021) $10^{5}$ Staicu et al. (Staicu and Pradel, 2018) 5s Davis et al. (Davis et al., 2018a) 10s Siddiq et al. (Siddiq et al., 2024a) 1s Shen et al. (Shen et al., 2018) $10^{8}$ Barlas et al. (Barlas et al., 2022) 1s Siddiq et al. (Siddiq et al., 2024b) 1s $10^{8}$ Davis et al. (Davis et al., 2019) 10s Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) 2s Wang et al. (Wang et al., 2023) $10^{5}$ McLaughlin et al. (McLaughlin et al., 2022) 10s Li et al. (Li et al., 2021) 1s Parolini et al. (Parolini and Miné, 2023) $10^{10}$ Wüstholz et al. (Wüstholz et al., 2017) Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) $\mu+3\sigma$ Li et al. (Li et al., 2022) 10s Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) 1s Davis et al. (Davis et al., 2021) Turoňová et al. (Turoňová et al., 2020) 10s, 100s
@@
Moreover, the inconsistency extends to the measurement of introduced slowdowns. Various papers use different metrics to quantify slowdowns, further complicating the landscape and hindering comparability across studies. Turoňová et al. (Turoňová et al., 2022) measure slowdowns in $10s$ and $100s$, Staicu et al. (Staicu and Pradel, 2018) in $5s$, Davis et al. (Davis et al., 2018a, 2019) in $10s$, Siddiq et al. (Siddiq et al., 2024a), Li et al. (Li et al., 2021) and Barlas et al. (Barlas et al., 2022) in $1s$, Bhuiyan et al. (Bhuiyan et al., 2023) in $2s$. Meanwhile, McLaughlin et al. (McLaughlin et al., 2022) use $10s$, and Bai et al. (Bai et al., 2021) define a problematic slowdown as the average response time ($\mu$) plus three times the standard deviation ($3\sigma$), to account for significant deviations from the norm. These excessively high thresholds likely miss many potential vulnerabilities that could cause significant issues under real-world conditions. In our experiments, we saw that under one-second slowdowns can be weaponized against a target server, with modest attacker resources (Bhuiyan and Staicu, 2022). Other papers rely on heuristics to define a vulnerability. For example, Davis et al. (Davis et al., 2017) utilize safe-regex (noa, 2024a), which uses static analysis rules to determine whether a regex is vulnerable. These tools are unsound and report both false positives and negatives.
@@ 3.4. Evaluations
Table 4. Summary of evaluation methods used in the reviewed studies. Most works (33/34) measure the success of a ReDoS attack from the perspective of the attacker, rather than from its impact on benign users (Quality of Service degradation). Over half of the works (25/34) are evaluated on a single language’s regex engine.
Paper Attack Simulation Measure QoS Degradation JavaScript Java Python Ruby PHP Go Perl Rust C# Su et al. (Su et al., 2024) Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) Liu et al. (Liu et al., 2021) Staicu et al. (Staicu and Pradel, 2018) Davis et al. (Davis et al., 2018a) Siddiq et al. (Siddiq et al., 2024a) Shen et al. (Shen et al., 2018) Barlas et al. (Barlas et al., 2022) Siddiq et al. (Siddiq et al., 2024b) Davis et al. (Davis et al., 2019) Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) Noller et al. (Noller et al., 2018) Wang et al. (Wang et al., 2023) Rathnayake et al. (Rathnayake and Thielecke, 2017) McLaughlin et al. (McLaughlin et al., 2022) Li et al. (Li et al., 2021) Parolini et al. (Parolini and Miné, 2023) Kirrage et al. (Kirrage et al., 2013) Wüstholz et al. (Wüstholz et al., 2017) Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) Li et al. (Li et al., 2022) Davis et al. (Davis, 2019) Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) Tandon et al. (Tandon et al., 2023) Weideman et al. (Weideman et al., 2016) Turoňová et al. (Turoňová et al., 2020)
@@
Language: Regular expression matching is comparatively slow in languages like JavaScript, Java, Python, and Ruby (Davis et al., 2019), leading to a significant amount of research focusing on detecting and defending ReDoS attacks in these languages. Reflecting this emphasis, 19 (54%) papers address JavaScript, 18 (51%) address Java, and 9 (25%) address Python, with JavaScript being particularly prominent.
@@ 4. ReDoS Mitigations in Regex Engines
In Section 3, we discuss many papers that measure ReDoS or propose defenses. They often assume that the matching of a regex engine can be modeled as a backtracking search (Section 3.2). However, it is unclear whether this assumption holds in practice, particularly in the context of modern regex engines.
@@
We answer these questions in two ways. First, a systematic analysis was conducted on regex engines in nine popular programming languages (Section 4.1). This analysis drew upon first-party sources such as source code, issue trackers, language documentation, blog posts from regex engine maintainers, and CVE reports. We also examined academic literature that describes regex engines. Second, it was measured whether super-linear regexes improve performance in the latest versions of these engines (Section 4.2).
@@ 4.1. Regex Engine Analysis
Language (Implementation) Nature of the Defense Deployment Year LOC to Implement On by Default? Current ReDoS-Safeness JavaScript (Node.js—V8 (Nod, [n. d.]; V8J, [n. d.])) A Non-backtracking Engine: Thompson-style (Thompson, 1968) 2020 1558 (10.6%) Partial Ruby (MRI/CRuby (Rub, [n. d.])) A Non-backtracking Engine: Memoization (Davis et al., 2021), Resource Capping: Time-based 2022 1100 (4.7%) Partial C# (.NET (Wha, [n. d.])) Resource Capping: Time-based; Brzozowski Derivatives Engine (Moseley et al., 2023) 2012, 2022 5417 (34.7%) Partial Java (OpenJDK (Ope, [n. d.])) Caching 2016 1712 (35.5%) Partial PHP (Zend Engine (PHP, [n. d.]; php, [n. d.])) Resource Capping: Counters Partial Perl (perl5 (Per, [n. d.])) Caching, Resource Capping: Counters Partial Rust (rustc (rus, [n. d.])) A Non-backtracking Engine: Thompson-style Safe Go (gc (Int, [n. d.])) A Non-backtracking Engine: Thompson-style Safe Python (CPython (Pyt, [n. d.])) Not Safe
Selection Criteria. Many modern programming languages have multiple implementations, runtime environments, engines, interpreters, or compilers that might use different regex engines. For instance, JavaScript has multiple engines for executing the source code, which can be used in different environments, such as V8 in Node.js and SpiderMonkey in Mozilla Firefox. To address this complexity, we applied the following criteria when selecting a specific regex engine for languages with multiple options:
  1. (1)
    We prioritized engines that are widely used in production environments and are likely to be targets of ReDoS attacks. This criterion ensures our analysis reflects real-world impact and relevance. As an illustrative example, Node.js is used as web server by 3.4% of all websites, and remarkably, it is used by 98.8% of all websites whose server-side programming language is known to be JavaScript (Usa, [n. d.]; Dis, [n. d.]). This makes V8 the most pertinent choice for our analysis. We focus on reference implementations for languages where comprehensive usage data is not readily available. Such implementations are typically mentioned in official documentation of programming languages and serve as the standard for the language (e.g., CPython for Python).
  2. (2)
    Additionally, we only consider implementations with publicly-accessible source code. This allows for a thorough examination of the implementation details and recent modifications to the engine. For instance, in the case of Java, we selected OpenJDK, an open-source platform, over Oracle’s JDK.
An overview of our regex engine selections for this study and their rationale can be found in Section D.1.
@@
Original Algorithm: The main JavaScript engine, V8, previously used a backtracking Spencer-style algorithm. It was highly optimized for common-case regexes (Gruber, 2017). It uses an explicit automaton representation, optimizing it and then rendering it as native machine code for execution (Corry et al., [n. d.]).
ReDoS Defense: In 2020, responding to ReDoS concerns, the V8 developers implemented an extra Thompson-style non-backtracking regex engine (Bidlingmaier, 2021). The new engine guarantees linear-time complexity for all supported regexes, but as a trade-off, it does not support E-regex features. For regexes that are not supported, the original engine serves as a fallback. This update was introduced in V8 version 8.8 (2020), but in Node.js, it remains an experimental feature that is not enabled by default and must be activated via a feature flag. Thus, some Node.js applications may not yet benefit from this feature. All E-regex features still use the backtracking engine.
Engineering Effort: The new engine is 1558 LOC. The original was 14741 LOC (excluding platform-specific assembly code). The effort was $\sim$10of the original.
@@ 4.1.2. Ruby (MRI/CRuby)
Original Algorithm: Ruby’s default regex engine is based on Onigmo (k.t, 2024), a fork of the Oniguruma (k.k, 2024) library, which is a Spencer-style backtracking regex engine.
ReDoS Defense: Until Ruby 3.2 (2022), Ruby’s port of Onigmo did not have built-in mitigations against ReDoS (issue 5820). In Ruby 3.2, the developers added two ReDoS defenses (rub, [n. d.]; Cha, [n. d.]). First, based on a community proposal, [^2] https://bugs.ruby-lang.org/issues/19104 Ruby implemented Davis et al.’s memoization technique for Spencer’s algorithm as a defense against ReDoS attacks (Davis et al., 2021). Interestingly, the Ruby developers used full memoization rather than selective memoization despite the higher space complexity. Ruby does not apply memoization when E-regex features are encountered, instead offering a timeout mechanism like C#. The memoization defense is on by default.
Engineering Effort: The relevant pull requests added a total of 1100 LOC, compared to the previous Ruby 3.1.6 engine’s footprint of 23384 LOC. The ratio is 5%.
@@
Original Algorithm: The first two regex engines in .NET used backtracking algorithms (Bac, 2023; Reg, 2021). One was a simple opcode interpreter. The other used an optimized machine code representation like V8’s (Bac, 2023; Toub, 2023; Com, 2021; Gutierrez, 2004).
ReDoS Defense: Similar to Ruby, in .NET v4.5 (2012), .NET introduced a timeout mechanism to prevent excessive backtracking (Wha, 2023). The timeout is checked after at most $O(n)$ steps, where $n$ is the length of the input, balancing fidelity with overhead (Toub, 2022; Bes, 2023). In .NET 7 (2022), Microsoft added a non-backtracking regex engine (Toub, 2022). This engine is based on Moseley et al.  (Moseley et al., 2023) and Saarikivi et al. (Saarikivi et al., 2019), with its core algorithm using Brzozowski derivatives and guaranteeing linear-time matches. However, both the timeout mechanism and the non-backtracking engine are opt-in features.
@@ 4.1.4. Java (OpenJDK)
Original Algorithm: OpenJDK Java’s previous regex engine implemented an NFA-based Spencer-style backtracking algorithm for regex matching (Pat, [n. d.]).
ReDoS Defense: As of Java 22 (2024), OpenJDK Java’s regex engine does not document defenses against ReDoS. Despite the absence of official documentation, in the OpenJDK source code repository and issue tracker, we identified that some performance updates related to ReDoS issues were introduced in Java 9 (2016). [^3] https://github.com/openjdk/jdk/commit/b45ea89 The OpenJDK maintainers introduced a bounded caching mechanism, improving problematic behavior for some ReDoS scenarios. This ReDoS defense is enabled by default and affects all OpenJDK Java versions $\geq$9.
@@ 4.1.5. Other ReDoS-Vulnerable Engines
PHP (Zend Engine), Perl (perl5), and Python (CPython) are also known to be using Spencer-style backtracking regex engines (Friedl, 2006; Davis et al., 2019), which are vulnerable to ReDoS attacks. According to Davis et al. (Davis et al., 2021), PHP (PHP Documentation Team, 2025) and Perl utilize counter-based caps to prevent catastrophic backtracking as an alternative to C# and Ruby’s time-based caps. Python has neither built-in engine optimizations against ReDoS vulnerabilities nor a native timeout-like mechanism for regex evaluations. We have not found any documented updates or defenses against ReDoS vulnerabilities for these languages since the analysis of Davis et al. (Davis et al., 2021) in 2021.

### v4

@@ 3.3. ReDoS Definitions
Table 3. Summary of ReDoS vulnerability definitions in reviewed studies. The literature uses inconsistent criteria, with over half (19/29) defining ReDoS based on slowdown. ◐ denotes incomplete details; ● indicates use of heuristics; – means the study did not use or report that criterion.
Paper Slowdown Number of Instructions Heuristics
Su et al. (Su et al., 2024) ¿10000
Davis et al. (Davis et al., 2017)
Turoňová et al. (Turoňová et al., 2022) 10s, 100s
Liu et al. (Liu et al., 2021) $10^{5}$
Staicu et al. (Staicu and Pradel, 2018) 5s
Davis et al. (Davis et al., 2018a) 10s
Siddiq et al. (Siddiq et al., 2024a) 1s
Shen et al. (Shen et al., 2018) $10^{8}$
Barlas et al. (Barlas et al., 2022) 1s
Siddiq et al. (Siddiq et al., 2024b) 1s $10^{8}$
Davis et al. (Davis et al., 2019) 10s
Kluban et al. (Kluban et al., 2022)
Bhuiyan et al. (Bhuiyan et al., 2023) 2s
Wang et al. (Wang et al., 2023) $10^{5}$
McLaughlin et al. (McLaughlin et al., 2022) 10s
Li et al. (Li et al., 2021a) 1s
Parolini et al. (Parolini and Miné, 2023) $10^{10}$
Wüstholz et al. (Wüstholz et al., 2017)
Petsios et al. (Petsios et al., 2017)
Kluban et al. (Kluban et al., 2024)
Demoulin et al. (Demoulin et al., 2019)
Chida et al. (Chida and Terauchi, 2022)
Bai et al. (Bai et al., 2021) $\mu+3\sigma$
Li et al. (Li et al., 2022) 10s
Li et al. (Li et al., 2021b)
Hassan et al. (Hassan et al., 2023)
Davis et al. (Davis et al., 2018b) 1s
Davis et al. (Davis et al., 2021)
Turoňová et al. (Turoňová et al., 2020) 10s, 100s
@@
Moreover, the inconsistency extends to the measurement of introduced slowdowns. Various papers use different metrics to quantify slowdowns, further complicating the landscape and hindering comparability across studies. Turoňová et al. (Turoňová et al., 2022) measure slowdowns in $10s$ and $100s$, Staicu et al. (Staicu and Pradel, 2018) in $5s$, Davis et al. (Davis et al., 2018a, 2019) in $10s$, Siddiq et al. (Siddiq et al., 2024a), Li et al. (Li et al., 2021a) and Barlas et al. (Barlas et al., 2022) in $1s$, Bhuiyan et al. (Bhuiyan et al., 2023) in $2s$. Meanwhile, McLaughlin et al. (McLaughlin et al., 2022) use $10s$, and Bai et al. (Bai et al., 2021) define a problematic slowdown as the average response time ($\mu$) plus three times the standard deviation ($3\sigma$), to account for significant deviations from the norm. These excessively high thresholds likely miss many potential vulnerabilities that could cause significant issues under real-world conditions. In our experiments, we saw that under one-second slowdowns can be weaponized against a target server, with modest attacker resources (Bhuiyan and Staicu, 2022). Other papers rely on heuristics to define a vulnerability. For example, Davis et al. (Davis et al., 2017) utilize safe-regex (Davis, 2024a), which uses static analysis rules to determine whether a regex is vulnerable. These tools are unsound and report both false positives and negatives.
@@ 3.4. Evaluations
Table 4. Summary of evaluation methods used in the reviewed studies. Most works (33/34) measure the success of a ReDoS attack from the perspective of the attacker, rather than from its impact on benign users (i.e., Quality of Service (QoS) degradation). ● indicates that the criterion or platform was explicitly considered; ◐ denotes incomplete details; – means the study did not use or report that criterion.
Paper Attack Simulation Measure QoS Degradation
JavaScript
Java
Python
Ruby
PHP
Go
Perl
Rust
C#
Su et al. (Su et al., 2024)
Davis et al. (Davis et al., 2017)
Turoňová et al. (Turoňová et al., 2022)
Liu et al. (Liu et al., 2021)
Staicu et al. (Staicu and Pradel, 2018)
Davis et al. (Davis et al., 2018a)
Siddiq et al. (Siddiq et al., 2024a)
Shen et al. (Shen et al., 2018)
Barlas et al. (Barlas et al., 2022)
Siddiq et al. (Siddiq et al., 2024b)
Davis et al. (Davis et al., 2019)
Kluban et al. (Kluban et al., 2022)
Bhuiyan et al. (Bhuiyan et al., 2023)
Noller et al. (Noller et al., 2018)
Wang et al. (Wang et al., 2023)
Rathnayake et al. (Rathnayake and Thielecke, 2017)
McLaughlin et al. (McLaughlin et al., 2022)
Li et al. (Li et al., 2021a)
Parolini et al. (Parolini and Miné, 2023)
Kirrage et al. (Kirrage et al., 2013)
Wüstholz et al. (Wüstholz et al., 2017)
Petsios et al. (Petsios et al., 2017)
Kluban et al. (Kluban et al., 2024)
Demoulin et al. (Demoulin et al., 2019)
Chida et al. (Chida and Terauchi, 2022)
Bai et al. (Bai et al., 2021)
Li et al. (Li et al., 2022)
Davis et al. (Davis, 2019)
Li et al. (Li et al., 2021b)
Hassan et al. (Hassan et al., 2023)
Davis et al. (Davis et al., 2018b)
Tandon et al. (Tandon et al., 2023)
Weideman et al. (Weideman et al., 2016)
Turoňová et al. (Turoňová et al., 2020)
@@
Language: Regex matching is comparatively slow in programming languages like JavaScript, Java, Python, and Ruby (Davis et al., 2019), leading to a significant amount of research focusing on detecting and defending ReDoS attacks in these languages. Reflecting this emphasis, 19 (54%) papers address JavaScript, 18 (51%) address Java, and 9 (25%) address Python, with JavaScript being particularly prominent.
@@ 4. ReDoS Mitigations in Regex Engines
In §3, we discuss many papers that measure ReDoS or propose defenses. They often assume that the matching of a regex engine can be modeled as a backtracking search (§3.2). However, it is unclear whether this assumption holds in practice, particularly in the context of modern regex engines.
@@
We answer these questions in two ways. First, a systematic analysis was conducted on regex engines in nine popular programming languages (§4.1). This analysis drew upon first-party sources such as source code, issue trackers, language documentation, blog posts from regex engine maintainers, and CVE reports. We also examined academic literature that describes regex engines. Second, we measured whether super-linear regexes improved performance in the latest versions of these engines (§4.2).
@@ 4.1. Regex Engine Analysis
Language (Implementation) Nature of the Defense Deployment Year LOC to Implement On by Default? Current ReDoS-Safeness
JavaScript (Node.js—V8 (Foundation, 2024b; Project, 2024a)) A Non-backtracking Engine: Thompson-style 2021 1558 (10.6%) Partial
Ruby (MRI/CRuby (Community, 2024b)) Memoization, Resource Capping: Time-based 2022 1100 (4.7%) Partial
C# (.NET (Microsoft, 2024)) Resource Capping: Time-based; Brzozowski Derivatives Engine 2012, 2022 5417 (34.7%) Partial
Java (OpenJDK (Oracle Corporation, 2024a)) Caching 2016 1712 (35.5%) Partial
PHP (Zend Engine (Group, 2024b, c)) Resource Capping: Counter-based Partial
Perl (perl5 (Perl.org, 2024a)) Caching, Resource Capping: Counter-based Partial
Rust (rustc (Foundation, 2024h)) A Non-backtracking Engine: Thompson-style Safe
Go (gc (Team, 2024c)) A Non-backtracking Engine: Thompson-style Safe
Python (CPython (Foundation, 2024d)) Not Safe
Selection Criteria: Many modern programming languages have multiple implementations, runtime environments, engines, interpreters, or compilers that might use different regex engines. For instance, JavaScript has multiple engines for executing the source code, which can be used in different environments, such as V8 in Node.js and SpiderMonkey in Mozilla Firefox. To address this complexity, we applied two key criteria when selecting a specific regex engine for languages with multiple options. First, we prioritized engines with substantial real-world adoption, as these are the most likely targets of ReDoS attacks in practice. Second, in cases where usage data is sparse or fragmented, we selected reference implementations that are officially endorsed and openly maintained, enabling thorough analysis of ReDoS defenses. An overview of our regex engine selections for this study and their rationale can be found in Appendix D.1.
@@
Original Algorithm: The main JavaScript engine, V8, previously used a backtracking Spencer-style algorithm. It was highly optimized for common-case regexes (Gruber, 2017). It uses an explicit automaton representation, optimizing it and then rendering it as native machine code for execution (Corry et al., 2009).
ReDoS Defense: In 2021, responding to ReDoS concerns, the V8 developers implemented an extra Thompson-style non-backtracking regex engine (Bidlingmaier, 2021). The new engine guarantees linear-time complexity for all supported regexes, but as a trade-off, it does not support some E-regex features. All regexes with unsupported E-regex features still have to use the backtracking engine. This update was introduced in V8 version 8.8, but in Node.js, it remains an experimental feature that is not enabled by default and must be activated via a feature flag. Thus, some Node.js applications may not yet benefit from the new non-backtracking regex engine.
Engineering Effort: The new engine is 1558 LOC. The original was 14741 LOC (excluding platform-specific assembly code). The effort was $\sim$10% of the original.
@@ 4.1.2. Ruby (MRI/CRuby)
Original Algorithm: Ruby’s default regex engine is based on Onigmo (Takata, 2024), a fork of the Oniguruma (Kosako, 2024) library, which is a Spencer-style backtracking regex engine.
ReDoS Defense: Until Ruby 3.2 (2022), Ruby’s port of Onigmo did not have built-in mitigations against ReDoS. In Ruby 3.2, the developers added two ReDoS defenses (Naruse, 2022; Shepelev, 2023). First, based on a community proposal, [^2] See https://bugs.ruby-lang.org/issues/19104. Ruby implemented Davis et al.’s (Davis et al., 2021) memoization technique for Spencer’s algorithm as a defense against ReDoS attacks. Interestingly, the Ruby developers used full memoization rather than selective memoization despite the higher space complexity. Ruby does not apply memoization when unsupported E-regex features are encountered, instead offering a timeout mechanism similar to C# as the second kind of defense. [^3] See https://bugs.ruby-lang.org/issues/17837. The memoization defense is on by default.
Engineering Effort: The relevant pull requests added a total of 1100 LOC, compared to the Ruby 3.1.6 engine’s footprint of 23384 LOC. The ratio is 5%.
@@
Original Algorithm: The first two regex engines in .NET used backtracking algorithms (Contributors, 2023a, 2021b). One was a simple opcode interpreter. The other used an optimized machine code representation like V8’s (Contributors, 2023a; Toub, 2023; Contributors, 2021a; Gutierrez, 2004).
ReDoS Defense: In .NET v4.5 (2012), .NET introduced a timeout mechanism to prevent excessive backtracking (Contributors, 2023c). The timeout is checked after at most $O(n)$ steps, where $n$ is the length of the input, balancing fidelity with overhead (Toub, 2022; Contributors, 2023b). In .NET 7 (2022), Microsoft added a non-backtracking regex engine (Toub, 2022). This engine is based on Moseley et al.  (Moseley et al., 2023) and Saarikivi et al. (Saarikivi et al., 2019), with its core algorithm using Brzozowski derivatives and guaranteeing linear-time matches for supported regexes. The new engine raises an error upon encountering unsupported E-regex features, so the developers must use the backtracking engine in such cases. However, both the timeout mechanism and the non-backtracking engine are opt-in features.
@@ 4.1.4. Java (OpenJDK)
Original Algorithm: OpenJDK Java’s previous regex engine implemented an NFA-based Spencer-style backtracking algorithm for regex matching (Oracle Corporation, 2024b).
ReDoS Defense: As of Java 22 (2024), OpenJDK Java’s regex engine does not document defenses against ReDoS. Despite the absence of official documentation, in the OpenJDK source code repository and issue tracker, we identified that some performance updates related to ReDoS issues were introduced in Java 9 (2016). [^4] See https://github.com/openjdk/jdk/commit/b45ea89. The OpenJDK maintainers introduced a bounded caching mechanism, improving problematic behavior for some ReDoS scenarios. This ReDoS defense is enabled by default and affects all OpenJDK Java versions $\geq$9.
@@ 4.1.5. Other ReDoS-Vulnerable Engines
PHP (Zend Engine), Perl (perl5), and Python (CPython) are also known to be using Spencer-style backtracking regex engines (Friedl, 2006; Davis et al., 2019), which are vulnerable to ReDoS attacks. According to Davis et al. (Davis et al., 2021), PHP (PHP Documentation Group, 2024) and Perl (Ludwig, 2009) utilize “counter-based caps” to prevent catastrophic backtracking as an alternative to C# and Ruby’s “time-based caps”. Python has neither built-in engine optimizations against ReDoS vulnerabilities nor a native timeout-like mechanism for regex evaluations. We have not found any new documented updates or defenses against ReDoS for these languages since the analysis of Davis et al. (Davis et al., 2021) in 2021.